# Heights and periods of algebraic cycles in families

### v1

@@ Acknowledgements
We would like to thank Richard Hain for numerous inspiring discussions and for sharing his notes and Xinyi Yuan for his help during the preparation of this project. Part of this work was done when ZG visited SZ at Princeton University in January, February, and May 2024. ZG would like to thank Princeton University for its hospitality. Finally, we would like to thank Richard Hain for mentioning our work and method at the conference “The Ceresa Cycle in Arithmetic and Geometry” at ICERM in May, 2024.
Ziyang Gao received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement n 945714). Shouwu Zhang is supported by an NSF award, DMS-2101787.
@@ 2. Period map associated with normal functions
The first goal of this section is to define the period map $\varphi_{\nu}$ associated with $\nu$ (when $\nu$ is clearly in context, we simply denote it by $\varphi$). This map fits into the diagram in the category of complex varieties
(2.1) ${{\mathscr{D}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{u}$${S\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\varphi}$${\Gamma\backslash{\mathscr{D}}}$
where
  1. (i)
@@
  2. (ii)
    $\Gamma$ is a suitable arithmetic subgroup of $\mathbf{G}({\mathbb{Q}})$.
We emphasize that $V$ is, in general, not a fiber of ${\mathbb{V}}_{{\mathbb{Q}}}$, and its geometric meaning will be given in Remark 2.2.
@@ 2.1. Universal period map to the classifying space
Recall that $\nu$ induces a VMHS $({\mathbb{E}}_{\nu},W_{\bullet},{\mathscr{F}}_{{\mathbb{E}}}^{\bullet})$ on $S$ of weight $-1$ and 00 which is graded-polarized (better, admissible), fitting into the short exact sequence in the category of graded-polarized VMHS $0\rightarrow{\mathbb{V}}_{{\mathbb{Z}}}\rightarrow{\mathbb{E}}_{\nu}\rightarrow{\mathbb{Z}}(0)_{S}\rightarrow 0$ with the canonical polarization on ${\mathbb{Z}}(0)_{S}$.
Let $u_{S}\colon\tilde{S}\rightarrow S^{\mathrm{an}}$ be the universal covering map. Then the pullback $u_{S}^{*}{\mathbb{E}}_{\nu}$ (resp. $u_{S}^{*}{\mathbb{V}}_{{\mathbb{Z}}}$) is canonically trivialized as a local system $u_{S}^{*}{\mathbb{E}}_{\nu}\cong\tilde{S}\times E_{\nu}$ (resp. $u_{S}^{*}{\mathbb{V}}_{{\mathbb{Z}}}\cong\tilde{S}\times M_{-1,{\mathbb{Z}}}$), with $E_{\nu}=H^{0}(\tilde{S},u_{S}^{*}{\mathbb{E}}_{\nu})$ (resp. with $M_{-1,{\mathbb{Z}}}=H^{0}(\tilde{S},u_{S}^{*}{\mathbb{V}}_{{\mathbb{Z}}})$). Then for each $s\in S({\mathbb{C}})$, the fiber ${\mathbb{E}}_{\nu,s}$ can be canonically identified with $E_{\nu}$ and ${\mathbb{V}}_{{\mathbb{Z}},s}$ can be canonically identified with $M_{-1,{\mathbb{Z}}}$. The pullback under $u_{S}$ of the short exact sequence of VMHS defined by $\nu$ becomes split after $\otimes{\mathbb{Q}}$ since $\tilde{S}$ is simply connected.
Each $\tilde{s}\in\tilde{S}$ gives rise to a polarized mixed Hodge structure $(E_{\nu,{\mathbb{Q}}},(W_{\bullet})_{\tilde{s}},({\mathscr{F}}_{{\mathbb{E}}}^{\bullet})_{\tilde{s}})$ on $E_{\nu,{\mathbb{Q}}}$, And this induces the universal period map
$\tilde{\varphi}=\tilde{\varphi}_{\nu}\colon\tilde{S}\longrightarrow{\mathscr{M}}$
@@
(2.2) ${{\mathscr{M}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{p}$${\tilde{S}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\tilde{\varphi}}$$\scriptstyle{\tilde{\varphi}_{0}}$${{\mathscr{M}}_{0}.}$
@@ 2.2. Construction of the Mumford–Tate domain
In practice, we need to refine this period map and replace the classifying space ${\mathscr{M}}$ by the smallest Mumford–Tate domain ${\mathscr{D}}$, which contains $\tilde{\varphi}(\tilde{S})$. It is constructed as follows.
@@
Fix $\tilde{s}\in\tilde{S}$. Define
(2.3) ${\mathscr{D}}:=\mathbf{G}({\mathbb{R}})^{+}\cdot\tilde{\varphi}(\tilde{s})\subseteq{\mathscr{M}}.$
@@
Finally, since $S$ is a quasi-projective variety, there exists an arithmetic subgroup $\Gamma$ of $\mathbf{G}({\mathbb{Q}})$ such that $\tilde{\varphi}$ descends to a morphism $\varphi\colon S^{\mathrm{an}}\rightarrow\Gamma\backslash{\mathscr{D}}$ fitting into the diagram (2.1).
@@ 2.4. Relating the Betti foliation and the period map
Recall the projection to the pure part $p\colon{\mathscr{D}}\rightarrow{\mathscr{D}}_{0}\subseteq{\mathscr{M}}_{0}$ from (2.2) and the semi-algebraic structure ${\mathscr{D}}=V({\mathbb{R}})\times{\mathscr{D}}_{0}$ as in Proposition B.3.(i). The diagram (2.1) can be complete into:
@@
For each $\tilde{s}\in\tilde{S}$, write $\tilde{s}_{V}$ for the image $\tilde{\varphi}(\tilde{s})$ under the natural projection ${\mathscr{D}}=V({\mathbb{R}})\times{\mathscr{D}}_{0}\rightarrow V({\mathbb{R}})$. This gives a map $\tilde{\varphi}_{V}\colon\tilde{S}\rightarrow V({\mathbb{R}})$ sending $\tilde{s}\mapsto\tilde{s}_{V}$. Notice that
(2.5) $\tilde{\varphi}=(\tilde{\varphi}_{V},\tilde{\varphi}_{0})\colon\tilde{S}\longrightarrow V({\mathbb{R}})\times{\mathscr{D}}_{0}={\mathscr{D}}.$
Recall $S_{{\mathscr{F}}}(t)=\{s\in S({\mathbb{C}}):\dim_{\nu(s)}(\nu(S)\cap{\mathscr{F}}_{\mathrm{Betti}})\geq t\}$ and define $S_{\mathrm{rk}}(t)=\{s\in S({\mathbb{C}}):\dim\nu_{\mathrm{Betti},s}(T_{s}S)\leq t\}$ for $\nu_{\mathrm{Betti},s}\colon T_{s}S\rightarrow T_{\nu(s)}{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}},s})$ defined in (1.3).
@@
$u_{S}^{-1}\left(S_{{\mathscr{F}}}(t)\right)$ =$=\{\tilde{s}\in\tilde{S}:\dim_{\tilde{s}}\tilde{\varphi}^{-1}(\{\tilde{s}_{V}\}\times{\mathscr{D}}_{0})\geq t\}$
(2.6) $=\bigcup\nolimits_{r\geq 0}\{\tilde{s}\in\tilde{S}:\dim_{\tilde{s}}\tilde{\varphi}^{-1}(\tilde{\varphi}(\tilde{s}))=r,~{}\{\tilde{s}_{V}\}\times\tilde{C}\subseteq\tilde{\varphi}(\tilde{S})\text{ for }$
@@
(2.7) $u_{S}^{-1}\left(S_{\mathrm{rk}}(t)\right)=\{\tilde{s}\in\tilde{S}:\mathrm{rank}(\mathrm{d}\tilde{\varphi}_{V})_{\tilde{s}}\leq t\}.$
Notice that in the union in (2.1), the second condition $\{\tilde{s}_{V}\}\times\tilde{C}\subseteq\tilde{\varphi}(\tilde{S})$ always holds true for $r>t$.
@@
Consider $\pi\colon{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})\rightarrow S$. Each $x\in{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$ lies in ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})_{s}={\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}},s})$ for $s=\pi(x)$, which is canonically isomorphic to $\mathrm{Ext}_{\mathrm{MHS}}({\mathbb{Z}}(0),{\mathbb{V}}_{{\mathbb{Z}},s})$ by Carlson [Car85]. Hence each $x$ gives rise to a ${\mathbb{Z}}$-mixed Hodge structure of weight $-1$ and 00, and for the universal covering map $u_{{\mathscr{J}}}\colon\tilde{{\mathscr{J}}}\rightarrow{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$ we obtain a period map $\tilde{\varphi}_{{\mathscr{J}}}\colon\tilde{{\mathscr{J}}}\rightarrow{\mathscr{M}}$. The map $\nu\circ u_{S}\colon\tilde{S}\rightarrow S\rightarrow{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$ lifts to $\tilde{\nu}\colon\tilde{S}\rightarrow\tilde{{\mathscr{J}}}$, and $\tilde{\varphi}=\tilde{\varphi}_{{\mathscr{J}}}\circ\tilde{\nu}$.
@@
The zero section of $\pi\colon{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})\rightarrow S$ gives rise to a Levi decomposition of $\mathbf{G}^{{\mathscr{M}}}$ and hence an identification ${\mathscr{M}}=M_{-1}({\mathbb{R}})\times{\mathscr{M}}_{0}$ as in Proposition B.3.(i) applied to ${\mathscr{M}}$, and $\tilde{\varphi}_{{\mathscr{J}}}$ becomes
(2.8) $\tilde{\varphi}_{{\mathscr{J}}}\colon\tilde{{\mathscr{J}}}=M_{-1}({\mathbb{R}})\times\tilde{S}\xrightarrow{(1,\tilde{\varphi}_{0})}M_{-1}({\mathbb{R}})\times{\mathscr{D}}_{0}\subseteq M_{-1}({\mathbb{R}})\times{\mathscr{M}}_{0}.$
By the last paragraph, the leaves of ${\mathscr{F}}_{\mathrm{Betti}}$ are precisely $u_{{\mathscr{J}}}\left(\tilde{\varphi}_{{\mathscr{J}}}^{-1}(\{a\}\times{\mathscr{M}}_{0})\right)$ for all $a\in M_{-1}({\mathbb{R}})$.
Thus $u_{{\mathscr{J}}}^{-1}\left(\nu(S_{{\mathscr{F}}}(t))\right)=\{\tilde{s}^{\prime}\in\tilde{\nu}(\tilde{S}):\dim_{\tilde{s}^{\prime}}\tilde{\varphi}_{{\mathscr{J}}}^{-1}(\{\tilde{s}^{\prime}_{V}\}\times{\mathscr{M}}_{0}\})\geq t\}$, where $\tilde{s}^{\prime}_{V}$ is the image of $\tilde{s}\in\tilde{{\mathscr{J}}}\xrightarrow{\tilde{\varphi}_{{\mathscr{J}}}}M_{-1}({\mathbb{R}})\times{\mathscr{M}}_{0}\rightarrow M_{-1}({\mathbb{R}})$ with the last map being the natural projection. Applying $\tilde{\nu}^{-1}$ (whose fibers are of dimension 00 because $\nu$ is injective) to the set above and noticing that $u_{{\mathscr{J}}}\circ\tilde{\nu}=\nu\circ u_{S}$, we have
$u_{S}^{-1}\left(S_{{\mathscr{F}}}(t)\right)=\{\tilde{s}\in\tilde{S}:\dim_{\tilde{s}}(\tilde{\varphi}_{{\mathscr{J}}}\circ\tilde{\nu})^{-1}(\{\tilde{s}^{\prime}_{V}\}\times{\mathscr{M}}_{0}\})\geq t\}.$
Hence the first equality in (2.1) holds true because $\tilde{\varphi}_{{\mathscr{J}}}\circ\tilde{\nu}=\tilde{\varphi}$ and by Lemma B.5 applied to ${\mathscr{D}}\subseteq{\mathscr{M}}$. Similarly, we get (2.7).
@@
The proof of Lemma 2.1 (with Lemma B.5 applied to ${\mathscr{D}}\subseteq{\mathscr{M}}$ and the identification $\tilde{{\mathscr{J}}}=M_{-1}({\mathbb{R}})\times\tilde{S}$) also yields the following assertion: ${\mathscr{J}}_{\nu}:=S\times_{\Gamma_{0}\backslash{\mathscr{D}}_{0}}(\Gamma\backslash{\mathscr{D}})\subseteq{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$ is the intermediate Jacobian of a sub-VHS translated by a torsion multisection; it contains $\nu(S)$ and is the minimal one containing $\nu(S)$ with respect to inclusion. The relative dimension $\dim{\mathscr{J}}_{\nu}-\dim S$ equals $\frac{1}{2}\dim V$.
@@
Since $\tilde{\varphi}_{V}$ factors through $\tilde{\varphi}$ and has target $V({\mathbb{R}})$, (2.7) immediately yields the following trivial upper bound
@@
3. The Betti rank and Zariski closedness of the rank-strata

### v2

@@ Acknowledgements
We would like to thank Ari Shnidman for asking us the question about the Northcott property for Ceresa cycles, Richard Hain for numerous inspiring discussions and for sharing his notes, Xinyi Yuan for his help during the preparation of this project, Dick Gross for suggesting we look at $\mathrm{Pic}^{1}({\mathscr{C}}_{g}/{\mathscr{M}}_{g})$ instead of just ${\mathscr{M}}_{g}$, and Peter Sarnak for asking the question about the positivity of the Beilinson–Bloch heights. Part of this work was completed when ZG visited SZ at Princeton University in January, February, and May 2024. ZG would like to thank Princeton University for its hospitality. We also appreciate your comments on an earlier draft.
Ziyang Gao received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (grant agreement n 945714). Shouwu Zhang is supported by an NSF award, DMS-2101787.
@@ 2.1. Universal period map to the classifying space
Recall that $\nu$ induces a VMHS $({\mathbb{E}}_{\nu},W_{\bullet},{\mathscr{F}}_{{\mathbb{E}}}^{\bullet})$ on $S$ of weight $-1$ and 00 which is graded-polarized (better, admissible), fitting into the short exact sequence in the category of graded-polarized VMHS
$0\rightarrow{\mathbb{V}}_{{\mathbb{Z}}}\rightarrow{\mathbb{E}}_{\nu}\rightarrow{\mathbb{Z}}(0)_{S}\rightarrow 0$
with the canonical polarization on ${\mathbb{Z}}(0)_{S}$.
Let $u_{S}\colon\widetilde{S}\rightarrow S^{\mathrm{an}}$ be the universal covering map. Then the pullback $u_{S}^{*}{\mathbb{E}}_{\nu}$ (resp. $u_{S}^{*}{\mathbb{V}}_{{\mathbb{Z}}}$) is canonically trivialized as a local system $u_{S}^{*}{\mathbb{E}}_{\nu}\simeq\widetilde{S}\times E_{\nu}$ (resp. $u_{S}^{*}{\mathbb{V}}_{{\mathbb{Z}}}\simeq\widetilde{S}\times V_{{\mathbb{Z}}}$), with $E_{\nu}=H^{0}(\widetilde{S},u_{S}^{*}{\mathbb{E}}_{\nu})$ (resp. with $V_{{\mathbb{Z}}}=H^{0}(\widetilde{S},u_{S}^{*}{\mathbb{V}}_{{\mathbb{Z}}})$). Then for each $s\in S({\mathbb{C}})$, the fiber ${\mathbb{E}}_{\nu,s}$ can be canonically identified with $E_{\nu}$ and ${\mathbb{V}}_{{\mathbb{Z}},s}$ can be canonically identified with $V_{{\mathbb{Z}}}$. The pullback under $u_{S}$ of the short exact sequence of VMHS defined by $\nu$ becomes split after $\otimes{\mathbb{Q}}$ since $\widetilde{S}$ is simply connected.
Each $\widetilde{s}\in\widetilde{S}$ gives rise to a polarized mixed Hodge structure $(E_{\nu,{\mathbb{Q}}},(W_{\bullet})_{\widetilde{s}},({\mathscr{F}}_{{\mathbb{E}}}^{\bullet})_{\widetilde{s}})$ on $E_{\nu,{\mathbb{Q}}}$, And this induces the universal period map.
$\widetilde{\varphi}=\widetilde{\varphi}_{\nu}\colon\widetilde{S}\longrightarrow{\mathscr{M}}$
@@
(2.1) ${{\mathscr{M}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{p}$${\widetilde{S}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\widetilde{\varphi}}$$\scriptstyle{\widetilde{\varphi}_{0}}$${{\mathscr{M}}_{0}.}$
In practice, we need to refine this period map $\widetilde{\varphi}$ and replace the classifying space ${\mathscr{M}}$ by the smallest Mumford–Tate domain ${\mathscr{D}}$ which contains $\widetilde{\varphi}(\widetilde{S})$. This construction is done in the next subsection.
@@ 2.2. Construction of the Mumford–Tate domain
Fix $\widetilde{s}\in\widetilde{S}$. Define
(2.2) ${\mathscr{D}}:=\mathbf{G}({\mathbb{R}})^{+}\cdot\widetilde{\varphi}(\widetilde{s})\subseteq{\mathscr{M}}.$
@@
Finally, since $S$ is a quasi-projective variety, there exists an arithmetic subgroup $\Gamma$ of $\mathbf{G}({\mathbb{Q}})$ such that $\widetilde{\varphi}$ descends to a morphism $\varphi\colon S^{\mathrm{an}}\rightarrow\Gamma\backslash{\mathscr{D}}$.
@@ 2.4. Relating the Betti foliation and the period map
Denote by ${\mathscr{D}}_{0}:=p({\mathscr{D}})\subseteq{\mathscr{M}}_{0}$ for the projection to the pure part $p$ from (2.1), and by abuse of notation $p=p|_{{\mathscr{D}}}\colon{\mathscr{D}}\rightarrow{\mathscr{D}}_{0}$. Recalling the semi-algebraic structure ${\mathscr{D}}=V({\mathbb{R}})\times{\mathscr{D}}_{0}$ as in Proposition B.3.(i), the period maps $\varphi$ and $\widetilde{\varphi}$ can be complete into:
@@
For each $\widetilde{s}\in\widetilde{S}$, write $\widetilde{s}_{V}$ for the image $\widetilde{\varphi}(\widetilde{s})$ under the natural projection ${\mathscr{D}}=V({\mathbb{R}})\times{\mathscr{D}}_{0}\rightarrow V({\mathbb{R}})$. This gives a map $\widetilde{\varphi}_{V}\colon\widetilde{S}\rightarrow V({\mathbb{R}})$ sending $\widetilde{s}\mapsto\widetilde{s}_{V}$. Notice that
(2.4) $\widetilde{\varphi}=(\widetilde{\varphi}_{V},\widetilde{\varphi}_{0})\colon\widetilde{S}\longrightarrow V({\mathbb{R}})\times{\mathscr{D}}_{0}={\mathscr{D}}.$
Recall $S^{\mathrm{Betti}}(t)=\{s\in S({\mathbb{C}}):\dim_{\nu(s)}(\nu(S)\cap{\mathscr{F}}_{\mathrm{Betti}})\geq t\}$ and define
$S_{\mathrm{rk}}(t)=\{s\in S({\mathbb{C}}):\dim\nu_{\mathrm{Betti},s}(T_{s}S)\leq t\}$
for $\nu_{\mathrm{Betti},s}\colon T_{s}S\rightarrow T_{\nu(s)}{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}},s})$ defined in (1.4).
@@
$u_{S}^{-1}\left(S^{\mathrm{Betti}}(t)\right)$ =$=\{\widetilde{s}\in\widetilde{S}:\dim_{\widetilde{s}}\widetilde{\varphi}^{-1}(\{\widetilde{s}_{V}\}\times{\mathscr{D}}_{0})\geq t\}$
(2.5) $=\bigcup\nolimits_{r\geq 0}\{\widetilde{s}\in\widetilde{S}:\dim_{\widetilde{s}}\widetilde{\varphi}^{-1}(\widetilde{\varphi}(\widetilde{s}))=r,~{}\{\widetilde{s}_{V}\}\times\widetilde{C}\subseteq\widetilde{\varphi}(\widetilde{S})\text{ for }$
@@
(2.6) $u_{S}^{-1}\left(S_{\mathrm{rk}}(t)\right)=\{\widetilde{s}\in\widetilde{S}:\mathrm{rank}(\mathrm{d}\widetilde{\varphi}_{V})_{\widetilde{s}}\leq t\}.$
Notice that in the union in (2.1), the second condition $\{\widetilde{s}_{V}\}\times\widetilde{C}\subseteq\widetilde{\varphi}(\widetilde{S})$ always holds for $r>t$.
@@
Consider $\pi\colon{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})\rightarrow S$. Each $x\in{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$ lies in ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})_{s}={\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}},s})$ for $s=\pi(x)$, which is canonically isomorphic to $\mathrm{Ext}_{\mathrm{MHS}}({\mathbb{Z}}(0),{\mathbb{V}}_{{\mathbb{Z}},s})$ by Carlson [Car85]. Hence each $x$ gives rise to a ${\mathbb{Z}}$-mixed Hodge structure of weight $-1$ and 00, and for the universal covering map $u_{{\mathscr{J}}}\colon\widetilde{{\mathscr{J}}}\rightarrow{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$ we obtain a period map $\widetilde{\varphi}_{{\mathscr{J}}}\colon\widetilde{{\mathscr{J}}}\rightarrow{\mathscr{M}}$. The map $\nu\circ u_{S}\colon\widetilde{S}\rightarrow S\rightarrow{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$ lifts to $\widetilde{\nu}\colon\widetilde{S}\rightarrow\widetilde{{\mathscr{J}}}$, and $\widetilde{\varphi}=\widetilde{\varphi}_{{\mathscr{J}}}\circ\widetilde{\nu}$.
@@
The zero section of $\pi\colon{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})\rightarrow S$ gives rise to a Levi decomposition of $\mathbf{G}^{{\mathscr{M}}}$ and hence an identification ${\mathscr{M}}=V({\mathbb{R}})\times{\mathscr{M}}_{0}$ as in Proposition B.3.(i) applied to ${\mathscr{M}}$, and $\widetilde{\varphi}_{{\mathscr{J}}}$ becomes
(2.7) $\widetilde{\varphi}_{{\mathscr{J}}}\colon\widetilde{{\mathscr{J}}}=V({\mathbb{R}})\times\widetilde{S}\xrightarrow{(1,\widetilde{\varphi}_{0})}V({\mathbb{R}})\times{\mathscr{D}}_{0}\subseteq V({\mathbb{R}})\times{\mathscr{M}}_{0}.$
By the last paragraph, the leaves of ${\mathscr{F}}_{\mathrm{Betti}}$ are precisely $u_{{\mathscr{J}}}\left(\widetilde{\varphi}_{{\mathscr{J}}}^{-1}(\{a\}\times{\mathscr{M}}_{0})\right)$ for all $a\in V({\mathbb{R}})$.
Thus $u_{{\mathscr{J}}}^{-1}\left(\nu(S^{\mathrm{Betti}}(t))\right)=\{\widetilde{s}^{\prime}\in\widetilde{\nu}(\widetilde{S}):\dim_{\widetilde{s}^{\prime}}\widetilde{\varphi}_{{\mathscr{J}}}^{-1}(\{\widetilde{s}^{\prime}_{V}\}\times{\mathscr{M}}_{0}\})\geq t\}$, where $\widetilde{s}^{\prime}_{V}$ is the image of $\widetilde{s}\in\widetilde{{\mathscr{J}}}\xrightarrow{\widetilde{\varphi}_{{\mathscr{J}}}}V({\mathbb{R}})\times{\mathscr{M}}_{0}\rightarrow V({\mathbb{R}})$ with the last map being the natural projection. Applying $\widetilde{\nu}^{-1}$ (whose fibers are of dimension 00 because $\nu$ is injective) to the set above and noticing that $u_{{\mathscr{J}}}\circ\widetilde{\nu}=\nu\circ u_{S}$, we have
$u_{S}^{-1}\left(S^{\mathrm{Betti}}(t)\right)=\{\widetilde{s}\in\widetilde{S}:\dim_{\widetilde{s}}(\widetilde{\varphi}_{{\mathscr{J}}}\circ\widetilde{\nu})^{-1}(\{\widetilde{s}^{\prime}_{V}\}\times{\mathscr{M}}_{0}\})\geq t\}.$
Hence the first equality in (2.1) holds because $\widetilde{\varphi}_{{\mathscr{J}}}\circ\widetilde{\nu}=\widetilde{\varphi}$ and by Lemma B.5 applied to ${\mathscr{D}}\subseteq{\mathscr{M}}$. Similarly, we get (2.6).
@@
The proof of Lemma 2.1 (with Lemma B.5 applied to ${\mathscr{D}}\subseteq{\mathscr{M}}$ and the identification $\widetilde{{\mathscr{J}}}=V({\mathbb{R}})\times\widetilde{S}$) also yields the following assertion: ${\mathscr{J}}_{\nu}:=S\times_{\Gamma_{0}\backslash{\mathscr{D}}_{0}}(\Gamma\backslash{\mathscr{D}})\subseteq{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$ is the intermediate Jacobian of a sub-VHS translated by a torsion multisection; it contains $\nu(S)$ and is the minimal one containing $\nu(S)$ with respect to inclusion. The relative dimension $\dim{\mathscr{J}}_{\nu}-\dim S$ equals $\frac{1}{2}\dim V$.
@@
Since $\widetilde{\varphi}_{V}$ factors through $\widetilde{\varphi}$ and has target $V({\mathbb{R}})$, (2.6) immediately yields the following trivial upper bound
@@
We can say more. By Griffiths’ transversality, we must have $\nu_{\mathrm{Betti},s}(T_{s}S)<\frac{1}{2}\dim V$ if ${\mathscr{J}}({\mathbb{V}}^{\prime}_{{\mathbb{Z}}})\rightarrow S$ is not an abelian scheme (or equivalently, the Hodge type of ${\mathbb{V}}^{\prime}_{{\mathbb{Z}}}\rightarrow S$ is not $(-1,0)+(0,-1)$); see above (D.3).
3. The Betti rank and Zariski closedness of the Betti strata